You are given a matrix 


"A=\left(\begin{array}{cccc}1 & -2 & -1\\ 2 & -3 & 0 \\1& -1 & -3\end{array} \right)."


Which of the following are true about A?

 

  1. "det (A)=-4"
  2. "adj (A)=\left(\begin{array}{cccc}9 & -5 & -3\\ 6 & -2 & -2 \\1& -1 & 1\end{array} \right)"
  3. "A^{-1}=\left(\begin{array}{cccc}-\frac{9}{4} & \frac{5}{4} & \frac{3}{4}\\ -\frac{3}{2} & \frac{1}{2} & \frac{1}{2} \\-\frac{1}{4}& \frac{1}{4} & -\frac{1}{4}\end{array} \right)"
  4. The system AX=0 has no solution, where 0 denotes a zero matrix of order 3×1.
  5. "adj (A)=\left(\begin{array}{cccc}9 & -5 & -3\\ 6 & -2 & -2 \\1& -1 & 1\end{array} \right)^{T}"
1
Expert's answer
2021-09-15T02:57:49-0400

"det(A)=\u22124"


"A^{-1}=\left(\begin{array}{cccc}-\frac{9}{4} & \frac{5}{4} & \frac{3}{4}\\ -\frac{3}{2} & \frac{1}{2} & \frac{1}{2} \\-\frac{1}{4}& \frac{1}{4} & -\frac{1}{4}\end{array} \right)"